Don Glover borrowed $13,000 for 250 days and paid $604.80 in simple interest on the loan. What annual simple interest rate did Don pay on the loan? (Round your answer to the nearest hundredth of a percent.)
Answer by fractalier(6550)

Simple interest is found by
I = Prt
If we wish the rate, we can find it by
r = I / Pt
Here we have
r = 604.80 / (13000)(250/365) = .068 = 6.8%